1) You own a portfolio that is invested as follows: $22,575 of Stock A, $3,750 of Stock B, $12,500 of Stock C, and $5,800 of Stock D. What is the portfolio weight of Stock B?

2) You own a $58,600 portfolio comprised of four stocks. The values of Stocks A, B, and C are $11,200, $17,400, and $20,400, respectively. What is the portfolio weight of Stock D?
Answer #1

1) Amount invested in stock A = 22575

Amount invested in stock B = 3750

Amount invested in stock C = 12500

Amount invested in stock D = 5800

Total amount invested = 22575+3750+12500+5800 = $44625

Weight of stock B = amount invested in stock B/Total amount invested = 3750/44625 = 0.0840336134453782 = 8.40336134453782%

Answer -> portfolio weight of stock B = 8.40336134453782%

2)Total amount invested in the Portfolio = VP = $58600

Value of stock A in the portfolio = VA = $11200

Value of stock B in the portfolio = VB = $17400

Value of stock C in the portfolio = VC = $20400

Value of stock D in the portfolio = VD

VA+VB+VC+VD = VP

11200+17400+20400 + VD = 58600

VD = 58600 - 49000 = 9600

Amount invested in Stock D = $9600

Portfolio weight of stock D = VD/VP = 9600/58600 = 0.16382252559727 = 16.382252559727%

Answer -> Portfolio weight of stock D = 16.382252559727%